OpenStudy (anonymous):

How many tablespoons are in 5 quarts?

OpenStudy (anonymous):

I think it is 320 16 tbs are in a cup, 4 cups are in a quart. 16 *4=64 *5=320

OpenStudy (yanasidlinskiy):

http://www.asknumbers.com/quarts-to-tablespoons.aspx

OpenStudy (anonymous):

its 320.